### v1

Abstract
Context. Robust impact monitoring of near-Earth objects is an essential task of planetary defense. Current systems such as NASA’s Sentry-II, University of Pisa’s CLOMON2 and Aegis of ESA have been highly successful, but independent approaches are essential to ensure reliability and to cross-validate predictions of possible impacts, probability and path on Earth.
Aims. We present NEOForCE (Near-Earth Objects’ Forecast of Collisional Events), a new independent monitoring system for asteroid impact prediction. By relying on orbital solutions from DynAstVO at Paris Observatory and using an original methodology for uncertainty propagation, NEOForCE provides an alternative line of verification for future impact assessments and strengthens the overall robustness of planetary defense.
Methods. As other monitoring system NEOForCE samples several thousands of so-called ’virtual asteroids’ from the uncertainty region and integrates their orbits up to 100 years into the future. Instead of searching for close approaches of the virtual asteroids themselves with the Earth, our system looks for times when the Earth comes close to the ”realistic” uncertainty regions around them, which are mostly stretched along their osculating orbits. For each virtual asteroid we also estimate the maximal impact probability, and only if this value is large enough do we continue to the next step. In this second step, we compute how the original asteroid orbit should be slightly modified so that the new trajectory leads to an Earth impact, which allows us to confirm the possible collision and estimate the impact probability.
Results. We tested NEOForCE against NASA’s Sentry-II system on five representative asteroids with high impact probability and significant number of possible collisions: 2000 SG344, 2005 QK76, 2008 JL3, 2023 DO, and 2025 JU. NEOForCE successfully recovered mostly all possible collisions reported by Sentry-II with impact probabilities above $10^{-7}$, demonstrating the robustness of our approach. In addition, NEOForCE identified several potential impacts at the $10^{-7}$$10^{-6}$ level that Sentry-II did not report.
Key Words.:
Celestial mechanics – Minor planets, asteroids: general – Methods: numerical
1 Introduction
Estimating the probability that a particular asteroid can hit the Earth in a foreseen future is an important task of planetary defense. The orbit of an asteroid in general does not have absolute or infinite precision, and for a newly discovered object the uncertainty of its orbital parameters can be significant. Moreover, this uncertainty increases with time if no additional observations are provided. Therefore, apart from the nominal solution of the orbital parameters (usually computed from the least-squares method), there are other solutions, which also well represent observations, sometimes with similar probability. These sets of orbital parameters are in general called virtual asteroids (VAs) (Milani et al., 2000). Some of these VAs can collide with the Earth and some pass by safely, which creates an impact probability different from zero or equal to one.
One of the most natural and general way to estimate the impact probability is the Monte-Carlo approach: sample a large number of VAs from the uncertainty region of orbital parameters according to the probability distribution, propagate their orbit in the future, and last derive impact probability by the fraction of colliding orbits over all all considered in the sample. This approach has many advantages: It does not have additional assumptions, and also has a direct formula for estimating the uncertainty of the obtained result: $\sigma_{MC}=\sqrt{\frac{P_{MC}(1-P_{MC})}{n}}$, where $P_{MC}$ is the impact probability and $n$ is the number of considered VAs (Vavilov & Medvedev, 2015). However, if the impact probability is small—which is generally the case with PHAs—the number of required VAs to integrate to roughly estimate the impact probability (so that $\sigma_{MC}\lesssim 0.3P_{MC}$) is $10/P_{MC}$. Therefore, this approach is impractical to systematic search for possible impacts with small impact probability values on a large set of of objects.
A much more effective approach is a linear approach: where one assumes that the errors of the orbital parameters at the time of a possible collision, $t$, are linearly related to the errors at the epoch of observations, $t_{0}$. In this case, one can integrate only the orbit of the nominal asteroid state vector together with variational equations, deriving the transition matrix and partial derivative matrix of orbital parameters at time $t$ over $t_{0}$. Using this matrix one can estimate the uncertainty region of the asteroid at time $t$ and assess the impact probability. The first linear method was introduced by Chodas (1993) who also proposed to use the target plane for the impact analysis. The important drawback of this approach is that the asteroid must come close to the Earth (enter its sphere of influence $\approx 929,000$ km), making linearization validity limited. Vavilov & Medvedev (2015) proposed the use of a special curvilinear coordinate system related to the asteroid’s osculating orbit to overcome this limitation. This allows us to calculate satisfactorily accurate probabilities of an asteroid collision with the Earth, even if the distance between the Earth and the nominal asteroid at time $t$ exceeds 1 au. Later Vavilov (2020) improved this approach and called it Partial Banana Mapping (PBM) method. Because the method constructs the curvilinear uncertainty region of an asteroid then finds the part of this region, which is closest to the Earth and maps it to the target plane. However, the linear relation between the errors of orbital parameters is a very important assumption and is limiting the usage of this method to the situation when the two-body formalism is not significantly violated.
A better compromise between linear methods and a full Monte-Carlo approach was introduced by Milani (1999) as a concept of line of variations (LOV). Instead of sampling the full multi-dimensional phase-space, the LOV reduces the problem to a single dimension that follows the direction where uncertainties and nonlinear effects are strongest (Milani et al., 2002, 2005a, 2005b). By distributing VAs along this line and interpolating between them, one can efficiently identify possible impacting solutions and trace distinct dynamical pathways. This approach makes it possible to detect even very small impact probabilities while keeping the computational effort manageable. Because of these advantages, operational monitoring systems at NASA, Sentry [^1] https://cneos.jpl.nasa.gov/sentry/, and CLOMON2222https://newton.spacedys.com/neodys2/index.php?pc=4.1 (operated by the University of Pisa) were both built upon the LOV technique. Recently, NASA introduced the Sentry-II system (Roa et al., 2021), which samples the full confidence region rather than restricting to LOV. Potential impacts are then confirmed by adding an impact pseudo-observation and refitting the orbit to match the data. These methods require several thousands of VAs’ orbits to integrate for the first search of close encounters only, but can detect possible impacts with probabilities of the order of $10^{-7}$ while being much more efficient than the Monte-Carlo one.
Vavilov & Hestroffer (2025) introduced a semi-linear method for impact probability estimation, as an extension of the linear Partial Banana Mapping method (Vavilov, 2020) built in previous work. This extension (called Impactor PBM or Improved PBM: IPBM) consists of a direct search for a state vector at the epoch of observations, $t_{0}$, that leads to a close encounter at the time of a possible collision, $t$. This method requires several orbits to be integrated, but still much less than required by Sentry of CLOMON2 systems. The method has a moderate limitation in the way that the asteroid position’s uncertainty should not exceed the whole orbit: if a VA that collides with the Earth makes different number of revolutions around the Sun than the nominal one, this collision can not be found. This naturally puts a threshold for the impact probability that can be more or less guaranteed to be found (order of $10^{-5}$). This method successfully fills the gap between restricted but very efficient linear methods and robust non-linear methods that require several thousands of orbit propagation.
In this work, we want to extend the applicability of this semi-linear method so that we can find possible collision with smaller impact probability values. We call our monitoring system Near-Earth Object’s Forecast of Collisional Events (NEOForCE). The idea is to take the approach of LOV and combine it with the semi-linear PBM. We sample VAs along LOV, assign smaller uncertainty regions for each VAs and apply semi-linear PBM for each VAs. The paper is organized as follows. In Introduction we briefly explain the concept of the PBM method. In Section 2 we describe step by step the algorithm of our system NEOForCE. Section 3 is dedicated to target plane analysis used in NEOForCE. In Section 4 we discuss the practical realization of the method, the completness of the system and validation by comparing with Sentry-II. This follows by a conclusion.
1.1 Concept of the Partial Banana Mapping method
We briefly outline the concept of the Partial Banana Mapping (PBM) method, which forms the core of our system. A more detailed discussion can be found in Vavilov & Medvedev (2015); Vavilov (2020)
The key assumption of PBM is that, over a considered time span, the uncertainties of the orbital elements evolve almost linearly from their values at the epoch of observations $t_{0}$. In other words, gravitational perturbations act in a similar way on the entire uncertainty region, so that most elements change coherently. In this case, the only orbital element that strongly differs is the mean anomaly —– the angle that determines the position of the asteroid along its orbit. As a result, the uncertainty region is stretched primarily along the orbital path, while remaining narrow and close to the nominal trajectory. This elongated, curved region resembles the shape of a banana (see Fig. 1).
Figure 1: Schematic illustration of the confidence curvilinear ellipsoid. Point A is the nominal position of the asteroid at time $t$, and point B is the VA on the main axis of the confidence ellipsoid at the same time $t$, which is closest to the Earth after projection onto its target plane. The arrow indicates the velocity direction of Earth with respect to the confidence ellipsoid. The bold line shows the nominal orbit of the asteroid. Image credit: Fig. 3 in (Vavilov, 2020).
To describe this ”banana-shaped” uncertainty region more accurately, we use a special coordinate system based on the osculating elliptical elements. In this system, one of the coordinates is the mean anomaly, $M$, and the other two $(\xi,\eta)$ describe small shifts to the side and above/below the orbit (Vavilov & Medvedev, 2015). This makes the uncertainty region easier to describe mathematically. In these curvilinear coordinates, the shape of the uncertainty region is often close to a six-dimensional Gaussian cloud.
The PBM method focuses on the part of the uncertainty region that might actually lead to an impact with the Earth. At the time of the possible impact it searches along the largest axis of the uncertainty region for the point (point B in Fig. 1) that comes closest to the Earth, after projecting everything onto a target plane (plane perpendicular to the asteroid-Earth relative motion). Once this closest point is found, PBM uses the local uncertainty shape around it to estimate the probability that the asteroid might hit the Earth. The method also finds the distance between point B and the Earth’s center in the target plane, and the difference in $M$ coordinates for the Earth and point A (nominal position of the asteroid).
2 NEOForCE monitoring system
Our impact monitoring system Near-Earth Object’s Forecast of Collisional Events uses a state vector (coordinates and velocities) and its 6-dimensional covariance matrix of an asteroid at the average time of observations. The covariance matrix $\mathbf{C}_{0}$ gives us an approximation of the uncertainty of the state vector $\mathbf{x}_{0}$ at time $t_{0}$. If inside this uncertainty region there is a vector of orbital parameters that leads to a collision with the Earth, then the asteroid could potentially have an impact. The goal of a monitoring system is to find all significantly large areas of orbital parameters that lead to collisions.
Our monitoring system consists of several stages. First, we cut the covariance matrix $\mathbf{C}_{0}$ by many pieces and take VAs as representatives of these pieces (Section 2.1 Initial sampling). Second, for each VAs we compute all the times when the collision can happen around this VA (Section 2.3 Close encounter precomputation). Third, we combine the situations from the previous step, which indicate to the same possible impact (Section 2.4 Selection of representatives). Forth, we try finding a Virtual Impactor (VI) or claim that this possible collision is spurious (Section 2.5 Search for virtual impactors). Finally, we combine similar VIs (Section 2.6 Combination of similar impactors).
2.1 Initial sampling
In general the uncertainty region of an asteroid, especially a newly discovered one, even at the mean epoch of observations is very elongated, in one preferred direction along an eigen vector. This preferred direction is also called Line of Variations (LOV) (Milani et al., 2002). On the first stage we sample virtual asteroids (VAs) from the uncertainty region along the LOV. This line is determined as the eigenvector of matrix $\mathbf{C}_{0}$ corresponding to the largest eigenvalue. Since matrix $\mathbf{C}_{0}$ is symmetric and positive defined, it can be decomposed by spectral decomposition:
$\mathbf{C}_{0}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\mathrm{T}}$ (1)
where $\mathbf{U}$ is an orthogonal matrix of the eigenvectors, and $\mathbf{\Lambda}$ is a diagonal matrix of the eigenvalues, $\mathrm{T}$ denotes the matrix transpose operation. The principal vector corresponding to the highest eigenvalue $\lambda_{11}$ determines the LOV.
We sample $n$ virtual asteroids (VAs) (in general $n=5001$) on this line according to Laplace distribution so that the Laplace integral ($\frac{\sqrt{2}}{2}\exp(-\sqrt{2}|x|)$) between two subsequent points is constant (and equals the integral from the last point to infinity). We use $\sqrt{2}$ to make the dispersion of the distribution equal one. Since we consider 5001 VAs the integral between the points equals $b=1/5002$. The VAs are then determined by the following:
$l_{i}=-\log(1-2ib)/\sqrt{2}*\sqrt{\lambda_{11}};\ i=-2500,2500,$ (2)
where $l_{i}$ is the difference along the LOV from the nominal state vector.
We consider each VA as a representative of its own vicinity. We assign the covariance matrix to each VA by exchanging the largest eigenvalue by $\lambda_{11}^{*}=\max(l_{i}-l_{i-1},l_{i+1}-l_{i})$. The covariance matrix determining the uncertainty region for a VA is computed by:
$\mathrm{C}^{*}_{0}=\mathbf{U}\mathbf{\Lambda}^{*}\mathbf{U}^{\mathrm{T}}\\$ (3)
where $\mathbf{\Lambda}^{*}$ is matrix $\mathbf{\Lambda}$ with $\lambda_{11}^{*}$ as the largest eigenvalue.
2.2 Initial check
Instead of running NEOForCE system for all near-Earth asteroids we first fulfill the initial check whether this asteroid possibly can come close to the Earth. We select 17 VAs along LOV (on 5-sigmasigma interval equidistantly), integrate their orbits into the future and compute minimal orbit intersection distance (MOID) between their orbits and the Earth’s orbit every 15 days. If the minimal MOID among all of them is greater than a certain threshold (0.002 au), we don’t run the computation of the impact probability at all.
2.3 Close encounter precomputation
On this step for each VA we compute the information about possible close encounters. We integrate the orbit of a VA with variational equations towards the future until the final time (in general, 100 years). On each integration step we compute the osculating orbit of a VA and then obtain the minimal distance between the position of the Earth and this osculating ellipse (or parabola/hyperbola) numerically. We then find the local minima in these distances indicating the time of the local minima $t$, the state vector of the VA, $\mathbf{x}^{*}_{t}$, and the Earth and the partial derivative matrix of VA’s state vector at $t$ over original time $t_{0}$, $\mathbf{\Phi}(t_{0},t)$. The covariance matrix of the VA for time $t$ is:
$\mathbf{C}^{*}_{t}=\mathbf{\Phi}(t_{0},t)\ \mathbf{C}^{*}_{0}\ \mathbf{\Phi^{\mathrm{T}}}(t_{0},t).$ (4)
If the minimal distance between the osculating orbit and the Earth is <0.05<0.05 au then we assume that a collision can not be ruled out at this time $t$ for a vicinity of this VA.
We then use the concept of the partial banana mapping method (see Sec. 1.1). We construct the banana-shaped uncertainty region at time $t$ for this VA using covariance matrix $\mathbf{C}^{*}_{t}$ restricting it to 5-$\sigma$ uncertainty. In this region we find the closest point to the Earth (point B in Fig. 1), its coordinates in $(\xi,\eta,M)$ coordinate system and compute the distance between the Earth and the projection of point B to its target plane. We also compute and save the difference in mean anomaly, $M$ coordinate, between the found point B and point A.
We also estimate the maximal impact probability for this VA assuming that its orbit leads to a collision with the Earth at time $t$, approximately. For this, we compute the full covariance matrix at time $t$ around this VA:
$\mathbf{\hat{C}}_{t}=\mathbf{\Phi}(t_{0},t)\ \mathbf{C}_{0}\ \mathbf{\Phi^{\mathrm{T}}}(t_{0},t).$ (5)
Then we hypothetically put the Earth at the center of the position of the VA at time $t$ (point A in Fig. 1) and compute the probability by the target plane method (Chodas, 1993). To account for the distance from the VA to the nominal orbit we then multiply this value by $\exp(-\frac{1}{2}\frac{\delta^{2}}{\lambda_{11}})$, where $\delta$ is the deviation in terms of LOV of this VA from zero (i.e. the distance on LOV of the considered VA from the nominal asteroid). This probability tells us what is the maximal possible value of the impact probability with the Earth selected asteroid can have at approximately time $t$ around the vicinity of this VA. If it is lower than our threshold (nominally $5\times 10^{-8}$) we don’t need to go to next stage to find virtual impactors. We consider this probability too low.
2.4 Selection of representatives
The information obtained in the previous step often indicates that several neighboring virtual asteroids (VAs) along the Line of Variation (LOV) may correspond to the same potential encounter with the Earth. In particular, if a given close-approach time $t$ is found, multiple VAs clustered around each other can all signal the possibility of a collision at that epoch. To avoid redundancy and to efficiently characterize each event, we identify a small set of representative VAs.
Figure 2: Schematic illustration of the different scenarios for grouping of VAs. The top chart shows the situation when the Earth is between two subsequent VAs (difference in $M$ has opposite signs). In this case we choose j+1-th VA as our primary representative and j-th as a secondary one. The bottom chart shows the situation when all the subsequent VAs are on one side from the Earth (difference in $M$ are all positive). In this situation we choose j+2-th as our primary representative. Note, that j+7-th VA is in another group but it will also be a representative of its group.
For each candidate epoch $t$, we first collect all VAs that point to a possible collision at approximately that time withing 10 days. We then divide them by groups so that each group consists of subsequent VAs along LOV. Within such a group we search for two scenarios:
  • Earth located between two consecutive VAs (upper chart in Fig. 2).
    If the difference in mean anomaly, $M$, computed in Section 2.4, shows opposite signs for two consecutive VAs, this implies that the Earth lies between them at the given time $t$. In this case, we select the VA that is closest to the Earth at approximately time $t$. This VA becomes the selected representative. Note that this situation can happen several times within one group if the spread of VAs exceeds several revolutions around the Sun. The second VA we consider as ’sidekick’.
  • Earth located on one side of the group (bottom chart in Fig. 2).
    If the Earth is not bracketed by two VAs (i.e., all $M$ differences have the same sign within the group), then we select only one VA — the one lying on the outer edge of the group and at the minimal distance to the Earth. This VA becomes the representative for the entire group.
Once we selected the representatives we apply additional filters to them. First, as was mentioned above, we consider only the ones with maximal impact probability $5\times 10^{-8}$. Second, the distance between the Earth and the curved uncertainty region (distance between point B and the Earth) for at least one of the VAs must be smaller than $0.002$ au. In the case where the Earth is located between two subsequent VAs even though we select only the closest to the Earth, the representative is still selected if these two conditions are fulfilled for one of the VAs. If the conditions are not met, we do not expect that the collision can happen or that the probability is significant. And third, we only consider the ones for which the distance between the Earth and the VA at the time of the possible collision $t$ smaller than $1$ astronomical unit. This limit with 1 au distance allows us to exclude the situations where the signs of $M$ coordinate differences are opposite but the actual positions of the VAs are extremely far from the Earth (on the different side of the osculating orbit).
This filter allows us to consider only the situation where we believe the collision can happen, and only significant ones with sufficient impact probabilities. After this preselection, the next stage will find explicit orbital parameters around each selected VA that can lead to a collision, and estimate the impact probability.
2.5 Search for virtual impactors
Again, we follow the idea of Milani et al. (2005a) in CLOMON2 system, that in order to confirm a possible collision, one must explicitly find a virtual impactor (a state vector at initial epoch, $t_{0}$, which actually leads to a collision). Therefore, in this stage we try to find the virtual impactor (VI) around each representative found in the previous step. For this we use a semi-linear approach extension of the PBM, briefly described here, for details see Vavilov & Hestroffer (2025).
The objective is to find the initial condition at epoch $t_{0}$ of the asteroid, bringing it at time $t$ to the closest point B in Fig. 1. The linear approximation reads as follows:
$\bm{\xi}_{*}-\bm{\xi}=\left[\frac{\partial\xi}{\partial x_{0}}\right](\bm{x}_{0*}-\bm{x}_{0}),$ (6)
where $\bm{\xi}$ and $\bm{\xi}_{*}$ are vectors in $(\xi,\eta,M)$ coordinate system at time $t$ of the asteroid (point A) and VA closest to the Earth on the main axis of curvilinear uncertainty ellipsoid (point B), correspondingly. $\bm{x}_{0}$ and $\bm{x}_{0*}$ are the state vectors of the selected VA and the one we are looking for, correspondingly, at $t_{0}$ and $\left[\frac{\partial\xi}{\partial x_{0}}\right]$ is the matrix of the partial derivatives of curvilinear coordinates at time $t$ over the Cartesian coordinates at $t_{0}$.
$(\bm{x}_{0*}-\bm{x}_{0})$ is the vector we are looking for: how to change the state vector of the selected VA on the previous stage so that after orbit integration it goes to point B at time $t$. In practice, for NEOForCE system we also consider: $\frac{3}{4}(\bm{x}_{0*}-\bm{x}_{0})$, $\frac{1}{2}(\bm{x}_{0*}-\bm{x}_{0})$ and state vectors. We integrate these orbits until time $t$ plus 0.5 days. If none of them collide with the Earth at approximately time $t$ we select the closest to the Earth. This newly selected VA is closer to the Earth at approximately time $t$ than the original one, so we repeat the procedure using this VA as the new reference (starting from computing $\bm{\xi}$, $\bm{\xi}_{*}$ vectors and matrix $\left[\frac{\partial\xi}{\partial x_{0}}\right]$). In this way, the method works as an iterative refinement process: at each step the candidate VA is moved closer to the Earth at $\approx t$ until one of three conditions is met: a collision is found, the solution drifts outside the 9-$\sigma$ uncertainty ellipsoid, or the maximum number of 10 iterations is reached.
In an extreme case where the stretching of the uncertainty region is very high and mismatches the linear estimation, the closest VA between these 3 cases can be further away than the original one. In this case we select $\frac{1}{6}(\bm{x}_{0*}-\bm{x}_{0})$ virtual asteroid and proceed again. Once the VA entered the sphere of influence (approximately 929,000 km) we launch the target plane analysis (see Section 3).
If virtual impactor is found, $\bm{x}_{0*}$, we estimate its impact probability by the target plane method and multiply this value by
$\exp{\left[-\frac{1}{2}[\bm{x}_{0*}-\bm{x_{0}}]^{\mathrm{T}}\mathbf{C}_{0}^{-1}[\bm{x}_{0*}-\bm{x_{0}}]\right]},$ (7)
which takes into account that the found $\bm{x}_{0*}$ is not the nominal one. If a VI is not found we can use a ’sidekick’ try to find a VI around it. We don’t do it by default in order to save the performance but we keep this as an option for special cases.
Figure 3: Schematic illustration of the work of NEOForCE system. On the upper chart one can see the nominal asteroid (purple dot), the original uncertainty region at the epoch of observation (black ellipse), the LOV (horizontal line), VAs sampled along LOV (black dots). The green area represents the smaller uncertainty region assigned to this VA. The red denotes the VI. On the bottom chart one can see the nominal position of the asteroid at time of a possible collision, $t$ (point A), the osculating orbit of the nominal asteroid (black part of an ellipse), the Earth, the position of the VA at time $t$ (green dot), the uncertainty region assigned for this VA at time $t$ (green curved ellipse). The red dot (point B) represents the position of the VI. The dashed vector shows the relative motion of the Earth and the green uncertainty region. The Earth collides with the red dot.
A schematic illustration of the work of our system is presented in Fig. 3.
2.6 Combination of similar impactors
On the last stage we check if some of the found VIs are essentially the same and point to the same possible collision. For this we integrate the orbits of each VI until the impact, saving their positions every 30 days. If at each integration step several VIs are close to each other (distance $10^{-3}$ au), we take only the one with the highest impact probability value. We publish a possible collision if the impact probability is higher than $10^{-10}$. The reason why this threshold is significantly higher than $5\cdot 10^{-8}$ threshold for the maximal impact probability is that if the maximal impact probability is too small it means the uncertainty of the asteroid around this area is too large. On the other hand, if the maximal impact probability is not too small but the actual impact probability of a VI is very low, this means that the uncertainty region only grazes the Earth rather than passing through its center. In such cases, even small changes in the nominal orbital solution can lead to large fluctuations in the estimated impact probability, which is why these cases are interesting.
3 Target plane analysis
3.1 Classical approach
The target plane (TP) analysis begins with the integration of the asteroid’s orbit until its encounter with the Earth. Let $(x_{0},y_{0},z_{0},\dot{x}_{0},\dot{y}_{0},\dot{z}_{0})$ be the state vector of the asteroid at the observational epoch $t_{0}$, derived from orbital fitting, and let $\mathbf{C}_{0}$ be its variance–covariance matrix. Integrating the orbit with variational equations (Battin, 1964) up to the time $t$ when the asteroid enters the Earth’s sphere of influence yields the state vector $(x,y,z,\dot{x},\dot{y},\dot{z})$ and the partial derivative matrix
$\mathbf{\Phi}(t_{0},t)=\left(\begin{array}[]{ccc}\frac{\partial x}{\partial x_{0}}&\cdots&\frac{\partial x}{\partial\dot{z}_{0}}\\ \vdots&\ddots&\vdots\\ \frac{\partial\dot{z}}{\partial x_{0}}&\cdots&\frac{\partial\dot{z}}{\partial\dot{z}_{0}}\\ \end{array}\right).$ (8)
The variance-covariance matrix at time $t$ is then
$\mathbf{C}_{xyz\dot{x}\dot{y}\dot{z}}=\mathbf{\Phi}(t_{0},t)\ \mathbf{C}_{0}\ \mathbf{\Phi^{\mathrm{T}}}(t_{0},t),$ (9)
where $\mathrm{T}$ denotes the matrix transpose operation. Matrix $\mathbf{C}_{xyz\dot{x}\dot{y}\dot{z}}$ provides an approximation of the uncertainty region of an asteroid at time $t$.
The asteroid’s relative velocity with respect to the Earth, $\bm{v}_{rel}$, is nearly the same for all virtual asteroids (VAs) in this region, meaning they share almost the same direction and speed. The target plane is defined as the plane perpendicular to $\bm{v}_{rel}$. The orientation of the coordinate axes on the plane is arbitrary; in our implementation we define them as
[…]
where $\bm{\xi},\bm{\eta},\bm{\zeta}$ are the axis of the new system, with $\bm{\xi}$ and $\bm{\zeta}$ define the coordinates on the target plane and $\bm{z}$ is the z-axis in a Cartesian coordinate system. The center of the system is the asteroid.
To estimate impact probabilities, the 6D uncertainty region is projected onto the target plane. An asteroid will collide with the Earth if the distance from the Earth’s center is smaller than its radius $R_{\oplus}$ (see Fig. 4). Because of gravitational focusing, however, the asteroid’s trajectory inside the Earth’s sphere of influence is hyperbolic, so the effective cross-section is larger. The minimum distance $q$ is given by
$q=\frac{b}{\sqrt{1+\frac{v_{s}^{2}}{v_{rel}^{2}}}},$ (11)
where $v_{s}\approx 11.186\ \mathrm{km\ s^{-1}}$ is the Earth’s surface escape velocity and $b$ is the impact parameter (the distance on the target plane between the projections of the Earth’s and asteroid’s centers).
Figure 4: Schematic illustration of the target plane method. The ellipsoid represents the uncertainty region of an asteroid when it enters the sphere of influence, the plane is the target plane, the circle on the plane is the projection of the Earth, and the ellipse on the plane is the projection of the ellipsoid. Image credit: (Vavilov, 2018).
Thus, a collision occurs when $b<R_{\oplus}\sqrt{1+v_{s}^{2}/v_{\infty}^{2}}$. The effective collision radius is therefore $R^{\prime}_{\oplus}=R_{\oplus}\sqrt{1+v_{s}^{2}/v_{\infty}^{2}}$.
The mapping from Cartesian state space (relative to the asteroid) to target-plane coordinates can be written as $\bm{u}=\mathbf{B}\bm{w}$ with a $2\times 6$ matrix $\mathbf{B}$ and $\bm{w}$ being the relative to the nominal asteroid state vector. The covariance matrix on the target plane is then $\mathbf{C}_{TP}=\mathbf{B}\mathbf{C}_{xyz\dot{x}\dot{y}\dot{z}}\mathbf{B}^{\mathrm{T}}$ (Rao, 1952). Placing the nominal solution at the origin, the impact probability reads
$P=\frac{1}{2\pi|\mathrm{det}\mathbf{C_{TP}}|^{\frac{1}{2}}}\int\limits_{S_{R^{\prime}_{\oplus}}}e^{-(\bm{u}^{\mathrm{T}}\mathbf{C_{TP}}^{-1}\bm{u})/2}\mathrm{d}\bm{u},$ (12)
where $S_{R^{\prime}_{\oplus}}$ is a circle with a radius $R^{\prime}_{\oplus}$.
3.2 Beyond classical approach
The classical target plane approach works well at typical encounter velocities. However, when the relative velocity is low, the encounter spans several days and the Earth’s motion can no longer be approximated as rectilinear. In such cases, the method can be inaccurate. This problem arises for asteroid 2000 SG344: for a possible impact on 18 September 2069, NEOForCE identified a virtual impactor (VI), but the classical target plane method estimated zero probability because the projected $(\xi,\zeta)$ coordinates of the Earth (reminder that in our system the asteroid is in the center) at entry into the Earth’s sphere of influence were too far from the center, despite an actual collision days later.
To address this, we adopted a modified approach similar to Milani et al. (2005a). We extend the integration until the epoch of closest approach and construct a ”modified target plane (MTP)” perpendicular to the relative velocity at that instant. If the asteroid collides with the Earth, the MTP is defined at the impact point. Using the relative position and velocity at closest approach, we compute the unperturbed velocity $v_{\infty}$ and derive the corresponding unperturbed target-plane coordinates $(\xi_{TP^{*}},\zeta_{TP^{*}})$ from the MTP coordinates $(\xi_{MTP},\zeta_{MTP})$:
[…]
where $E$ is the total orbital energy, $\mu=GM_{\oplus}/R_{\oplus}$, and $M_{\oplus},R_{\oplus}$ are the Earth’s mass and radius. For high-speed encounters, $(\xi_{TP^{*}},\zeta_{TP^{*}})\approx(\xi_{TP},\zeta_{TP})$, but at low velocities the difference is significant.
Since the covariance matrix gives adequate estimation of the uncertainty region only before the entry into the sphere of influence, we approximate the mapping between TP1 (classical TP at entry) and TP2 (unperturbed TP built from MTP) as a rotation by an angle $\alpha$ and a scale factor $sc$, estimated from
$\begin{array}[]{l}sc=\sqrt{\frac{\xi_{TP^{*}}^{2}+\zeta_{TP^{*}}^{2}}{\xi_{TP}^{2}+\zeta_{TP}^{2}}}\\ \alpha=\mathrm{atan2}(\zeta_{TP^{*}},\xi_{TP^{*}})-\mathrm{atan2}(\zeta_{TP},\xi_{TP}).\end{array}$ (14)
For 2000 SG344, $sc$ can exceed 1.3 while $\alpha$ is only a couple of degrees. The relative velocity $\bm{v}_{\infty}$ is also significantly different from $\bm{v}_{rel}$, and hence the corrected gravitational focusing radius of the Earth, $R_{\oplus}^{\prime}$ changes. That is why we sometimes had a situation where we found a VI but the $(\xi_{TP},\zeta_{TP})$ coordinates are further away than the corrected radius of the Earth, which of course led to wrong impact probability values.
We therefore map $\mathbf{C}_{xyz\dot{x}\dot{y}\dot{z}}$ onto TP1, rotate by $\alpha$, and scale by $sc$ to obtain the $2\times 2$ matrix $\mathbf{C}_{TP^{*}}$.
3.2.1 Search for virtual impactor
To confirm whether a potential encounter corresponds to a genuine collision, we follow the method of Milani et al. (2005a). The 6D state vector at $t$ relative to the nominal state vector of the asteroid can be partitioned into $\bm{r},\bm{s}]$, where $\bm{r}=(\xi_{TP},\zeta_{TP})$ are target-plane coordinates and $\bm{s}$ is the complementary 4D vector.
We seek the adjusted state $\hat{\bm{r}},\hat{\bm{s}}$] corresponding to a collisional orbit with the highest probability density. The procedure is: 1. Compute $\bm{r}_{*}=(\xi_{TP^{*}},\zeta_{TP^{*}})$ from Eq. (13). 2. Obtain $\mathbf{C}_{TP^{*}}$ by rotating and scaling $\mathbf{C}_{TP}$. 3. Select $\hat{\bm{r}}_{*}$ inside $R^{\prime}_{\oplus}$ that minimizes $\hat{\bm{r}}_{*}^{\mathrm{T}}\mathbf{C}_{TP^{*}}^{-1}\hat{\bm{r}}_{*}$. 4. Transform back to TP1 to get $\hat{\bm{r}}$.
The covariance matrix $\mathbf{C}$ of $\bm{r},\bm{s}]$ at time $t$ can be decomposed as
$\mathbf{C}=\left(\begin{array}[]{cc}\mathbf{C_{r}}&\mathbf{C_{rs}}\\ \mathbf{C_{sr}}&\mathbf{C_{s}}\end{array}\right).$ (15)
According to (Milani et al., 2005a) , vector $\bm{s_{*}}$ can be found from
$\bm{s_{*}}=\bm{s}+\mathbf{C_{r}}^{-1}\mathbf{C_{sr}}(\bm{r}-\bm{r_{*}}).$ (16)
We then search for the the vector $\bm{x}_{0*}$ at epoch of observations, $t_{0}$, that leads to vector $\bm{r_{*}},\bm{s_{*}}]$ at time $t$, and hence, to the collision. The linear relation reads
$(\bm{r_{*}},\bm{s_{*}})]-[(\bm{r},\bm{s})]=\mathbf{W}\cdot\mathbf{\Phi}(t_{0},t)\ \ (\bm{\tilde{x}}_{0*}-\bm{x}_{0*}),$ (17)
where $\mathbf{W}$ is the $6\times 6$ rotation matrix into the TP2 frame. As in Section 2.5 we also consider $\frac{3}{4}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$ and $\frac{1}{2}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$ and select the closest one to the Earth at time $t$. If all these three are more distant than $\bm{x}_{0*}$ at time $t$ we take $\frac{1}{6}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$.
If the state vector $\bm{\tilde{x}}_{0*}$ does not lead to a collision approximately at time $t$, this process needs to be iterated taking $\bm{\tilde{x}}_{0*}$ as $\bm{{x}}_{0*}$ and using Eqs. (16) and (17). If the procedure does not determine a collision after 10 iterations or it is outside of 9-$\sigma$ 6D ellipsoid, we consider this impactor as spurious.
4 Practical realization and discussion
In this version of the system we used Everhart (1974) integrator of 15-th order. In the model we included gravitational influence of 8 major planets (and Pluto for historical reasons). We also considered the Earth and the Moon separately. The orbits of asteroids and their covariance matrices are taken from DynAstVO database [^3] https://epn.imcce.fr/dynastvo.html Desmars et al. (2017).
4.1 Completeness problem
The completeness of our system is determined by the threshold of impact probability that can be detected certainly. Following (Milani et al., 2005a) one can estimate the generic completion level: the highest impact probability that can be missed by our monitoring system if the uncertainty trail on the target plane is fully linear.
Under these assumptions, a possible collision can be missed if both representatives found in Section 2.4 are further than 1 au from the Earth. This literally means that the distance between two subsequent VAs at the time of this possible collision exceeds 2 au along the orbit. According to the sampling of VAs along the LOV, the maximal integral probability of a Gaussian distribution on the interval between two subsequent VAs is approximately $3\times 10^{-4}$. This is slightly higher than if we sampled VAs along the LOV according to Gaussian distribution. In this case, the probability would be $2\cdot 10^{-4}$. Our objective is hence the following: If we took a sample of 5001 VAs according to the Gaussian distribution, the latest VAs would be only 3.54 sigmas away from the nominal state vector ($l_{2500}=3.54$ from equation 2). By sampling the Laplace distribution $l_{2500}=5.53$ instead, we have VAs much further from the nominal state vector, without loosing much in terms of completeness. Hence, the highest impact probability that can be missed is approximately:
$\frac{D_{\oplus}[au]}{2\penalty 10000\ \mathrm{au}}3\times 10^{-4}$ (18)
To take into account the gravitational focusing of the Earth in order to properly estimate the completeness we need to increase the diameter of the Earth $D_{\oplus}$ by $\sqrt{1+v_{e}^{2}/v_{\infty}^{2}}$ (Chodas, 1993), where $v_{e}$ is the escape velocity from the Earth’s surface ($\approx 11.186$ km/s) and $v_{\infty}$ is the unperturbed geocentric velocity (approximately the velocity when crossing the sphere of influence). When the relative velocity of an asteroid and the Earth is high, the diameter of the Earth doesn’t change much and is about $10^{-4}$, which makes the completeness to be around $1.5\times 10^{-8}$. However, for a very slow moving asteroids (with 1 km/s relative velocity) the diameter is 11.22 times larger, that lowers the completeness to $1.7\times 10^{-7}$. In Section 2.4 we exclude all the representatives with estimated maximal impact probability lower than $5\times 10^{-8}$, which sounds reasonable.
4.2 Comparison to Sentry
To evaluate the robustness of our system, we performed detailed comparisons using five representative asteroids: 2005 QK76, 2008 JL3, 2023 DO, 2025 JU, and 2000 SG344. We selected these asteroids from the web-page Sentry of Central for Near-Earth Objects Study (CNEOS) [^4] https://cneos.jpl.nasa.gov/sentry. We were looking for objects at the top of the list with high cumulative impact probability and significant number of possible impacts with probabilities higher than $5\times 10^{-7}$ in order to be able to make a comparison. The asteroids have different observational arcs (from 1.7 days to 1.4 years), and different number of observations (from 14 to 52). For each case, we adopted the same orbital solutions as those used by NASA’s Sentry-II system, retrieved via the NASA SSD API [^5] https://ssd-api.jpl.nasa.gov on September 9, 2025, so that the results would be directly comparable.
The outcomes of our system and Sentry-II are presented in the tables below. In each table, “Date” refers to the epoch of the possible collision identified by NEOForCE, while $P_{{NEOForCE}}$ and $P_{{NASA}}$ are the corresponding impact probabilities computed by our system and Sentry-II, respectively. We attempted to match virtual impactors (VIs) such that, for collisions occurring at similar epochs, both the probability and the distance metric $\sigma_{VI}$ are comparable.
In our approach, $\sigma_{VI}$ quantifies how far a VI lies from the nominal orbit in terms of standard deviations and is defined as
$\sigma_{VI}^{2}=[\bm{x}_{0*}-\bm{x_{0}}]^{\mathrm{T}}\mathbf{C}_{0}^{-1}[\bm{x}_{0*}-\bm{x_{0}}],$ (19)
where $\bm{x}_{0*}$ is the state vector of the VI, $\bm{x}_{0}$ is the nominal state vector, and $\mathbf{C}_{0}$ is the covariance matrix. Sentry-II estimates $\sigma{VI}$ using the root-mean-square residuals of the observational fit for the impacting solution.
It is important to note that similar $\sigma_{VI}$ values do not necessarily imply that the same VI was found: many virtual asteroids can lie on the same 5D ellipsoidal surface defined by a given $\sigma_{VI}$. Conversely, two very similar VIs may have different $\sigma_{VI}$ values, for example, a small shift along the direction of the smallest eigenvector of $\mathbf{C}_{0}$ can leave the overall trajectory nearly unchanged while significantly altering $\sigma_{VI}$.
Table 1: Comparison table for asteroid 2025 JU.
Date $P_{NEOForCE}$ $\sigma_{VI}$ $P_{NASA}$ $\sigma_{VI_{NASA}}$
2074-10-2.3 4.10e-05 1.091 4.30e-05 1.539
2074-10-2.3 1.49e-05 0.998 1.50e-05 1.500
2076-10-1.9 1.71e-07 3.268 1.70e-07 3.855
2079-10-2.5 1.91e-08 0.949 2.00e-08 1.362
2079-10-2.5 8.04e-08 0.947 8.00e-08 1.193
2079-10-2.5 6.25e-07 0.742 6.80e-07 0.956
2079-10-2.5 5.68e-07 0.741 5.50e-07 0.937
2091-10-2.4 5.50e-10 1.182
2091-10-2.5 3.56e-09 1.830
2103-10-3.3 5.25e-09 3.011
2108-10-2.6 2.21e-08 1.164 2.90e-08 1.929
2108-10-2.5 7.23e-08 1.170 6.70e-08 1.471
2108-10-2.6 3.13e-07 1.175 3.10e-07 1.675
4.2.1 Asteroid 2025 JU
Asteroid 2025 JU is newly discovered, with 38 observations spanned over 35 days. Its cumulative impact probability is about $6\times 10^{-5}$, not very high and it also has only a few significant possible impacts. Table 1 shows an iconic agreement between two systems with Sentry loosing three possible collisions with a probability less than $5.3\times 10^{-9}$ (way below the mentioned completeness level).
Table 2: Comparison table for asteroid 2005 QK76.
Date $P_{NEOForCE}$ $\sigma_{VI}$ $P_{NASA}$ $\sigma_{VI_{NASA}}$
2030-02-26.3 4.59e-05 0.156 3.80e-05 0.263
2038-02-26.3 1.63e-05 1.179 1.60e-05 2.314
2038-02-26.3 1.20e-08 2.556 7.10e-09 5.534
2038-02-26.2 1.71e-07 1.742
2041-02-26.0 6.40e-08 0.707
2041-02-26.0 1.66e-08 0.706
2048-02-26.8 3.41e-06 0.662 3.40e-06 3.625
2048-02-26.8 1.33e-05 0.604 1.30e-05 3.786
2059-02-26.5 5.58e-10 4.127 1.60e-10 6.973
2062-02-26.2 3.97e-08 2.220
2078-02-26.2 2.24e-10 2.924
2083-02-26.5 4.49e-08 3.123
4.2.2 Asteroid 2005 QK76
Asteroid 2005 QK76 has been discovered on August 30 2025. It has only 14 observations spanned over 1.67 days. Such a short observation arc makes its orbit highly uncertain and a good test for comparison. Sentry-II does not use the covariance matrix only the observations itself, while NEOForCE uses the covariance matrix in coordinates and velocities at the mean epoch of observations. As one can see from Table 2, there is no significant difference in the impact probabilities of the possible collisions found. However, the values of $\sigma_{VI}$ are substantially different, with larger values for Sentry-II. Sentry-II did not find the possible collision on February 26, 2038 with impact probability of $1.7\times 10^{-7}$, but this is at the completeness limit of Sentry-II.
Table 3: Comparison table for asteroid 2023 DO.
Date $P_{NEOForCE}$ $\sigma_{VI}$ $P_{NASA}$ $\sigma_{VI_{NASA}}$
2057-03-23.8 4.38e-04 0.690 4.40e-04 0.693
2058-03-24.2 4.47e-07 2.549 4.80e-07 2.555
2058-03-23.8 1.51e-08 0.689
2059-03-24.0 8.90e-06 0.714 8.40e-06 0.717
2060-03-23.3 5.12e-07 0.692
2061-03-24.0 8.42e-08 2.244
2061-03-23.9 3.92e-09 2.186
2061-03-23.8 6.65e-09 2.137
2062-03-24.0 8.05e-07 2.797 6.90e-07 2.803
2062-03-23.8 2.05e-07 0.683 2.10e-07 0.687
2063-03-24.4 6.52e-08 2.242
2064-03-23.3 8.18e-08 0.696 7.90e-08 0.706
2064-03-23.4 6.05e-07 0.679 5.90e-07 0.680
2064-03-23.6 3.50e-08 1.413
2065-03-23.9 2.59e-08 2.181 7.30e-08 2.245
2065-03-23.9 4.75e-06 1.780 4.80e-06 1.785
2066-03-24.1 2.12e-08 1.412 2.00e-08 1.417
2066-03-23.8 2.05e-07 0.697 1.90e-07 0.702
2066-03-23.9 1.25e-06 0.673 1.10e-06 0.674
2066-03-24.05 1.50e-09 2.388
2067-03-24.3 2.61e-08 2.192 2.00e-08 2.553
2067-03-24.4 1.40e-07 2.149 1.30e-07 2.154
2067-03-24.4 8.86e-08 2.292
2067-03-24.4 6.91e-08 2.180
2068-03-23.4 7.52e-07 0.664 8.60e-07 0.670
2068-03-23.1 9.42e-09 0.713
2069-03-23.8 7.34e-08 2.128 7.00e-08 2.133
2069-03-23.5 2.13e-09 0.679
2070-03-23.6 1.30e-08 0.713
2071-03-24.2 3.77e-09 2.795
2072-03-23.6 6.97e-08 2.410 6.80e-08 2.416
2072-03-23.5 3.71e-08 2.452
2072-03-23.5 4.04e-10 2.257
2072-03-23.5 2.43e-10 2.190
2072-03-23.5 1.29e-09 1.422
2072-03-23.5 8.30e-09 1.413
2072-03-22.9 7.58e-09 0.713
2073-03-23.8 4.47e-07 2.666 2.90e-07 2.671
2074-03-24.1 9.31e-09 2.408
2074-03-24.1 1.03e-09 2.383
2074-03-24.0 9.12e-09 1.838
2076-03-23.6 1.62e-10 2.458
2076-03-23.7 8.15e-09 2.248
2076-03-23.7 3.38e-08 1.418
2077-03-23.4 3.40e-08 0.664 3.40e-08 0.667
2077-03-23.3 1.69e-09 0.714
2078-03-23.9 4.39e-09 1.422 7.90e-09 2.306
2080-03-23.44 3.30e-08 1.841
2080-03-23.57 1.20e-08 2.154
2082-03-24.20 6.70e-08 2.252
2085-03-23.6 1.79e-06 2.221
2092-03-23.90 2.70e-09 1.385
4.2.3 Asteroid 2023 DO
This asteroid was observed over a relatively longer arc of 41 days, with 52 recorded observations. Discovered on February 13, 2023, it has 25 possible collisions published by Sentry-II. As shown in Table 3, our NEOForCE system identified 47 possible collisions in total. Of these, NEOForCE did not reproduce 5 solutions listed by Sentry-II, all with impact probabilities below $6.7\times 10^{-8}$, close to our completeness threshold of $5\times 10^{-8}$. Conversely, Sentry-II missed 27 possible collisions detected by NEOForCE, including one with a probability of $1.79\times 10^{-6}$ on March 23, 2085, another at $5.1\times 10^{-7}$ on March 23, 2060, and one at $8.4\times 10^{-8}$ on March 24, 2061.
4.2.4 Asteroid 2008 JL3
This asteroid was observed over a short arc of only 7 days, yielding 35 measurements in total. Discovered on May 2, 2008, it has 44 possible collision solutions listed in Sentry-II. A comparison of our system with Sentry-II (Table NEOForCE: Near-Earth Objects’ Forecast of Collisional Events) shows that NEOForCE identified 87 possible collisions but did not recover 3 mentioned by Sentry-II, all with impact probabilities below $4.6\times 10^{-9}$ — way below our completeness threshold. In contrast, Sentry-II missed 46 possible collisions, including one with the highest reported impact probability of $1.26\times 10^{-6}$ on April 28, 2109, as well as three others with probabilities exceeding $10^{-7}$.
4.2.5 Asteroid 2000 SG344
Table 4: Comparison table for asteroid 2000 SG344 (beginning).
Date $P_{NEOForCE}$ $\sigma_{VI}$ $P_{NASA}$ $\sigma_{VI_{NASA}}$
2069-09-18.5 1.96e-07 4.008 2.20e-07 3.995
2070-09-17.3 2.18e-04 1.732 2.30e-04 1.727
2071-09-16.0 1.03e-03 0.864 1.00e-03 0.862
2071-09-10.4 1.68e-04 0.720 1.40e-04 0.717
2072-02-3.5 3.59e-06 1.838 2.70e-06 1.832
2072-09-13.8 1.59e-06 3.860 1.40e-06 3.845
2072-09-8.5 5.06e-07 3.683
2073-02-1.7 1.44e-05 0.737 1.30e-05 0.735
2073-02-12.0 1.04e-05 1.790 1.00e-05 1.782
2073-02-9.2 2.06e-06 1.834 5.00e-06 1.827
2073-01-29.26 6.70e-07 1.839
2074-02-10.5 4.02e-05 0.787 4.10e-05 0.785
2074-02-8.5 4.48e-05 0.748 4.00e-05 0.745
2074-01-28.5 4.43e-06 0.739 3.40e-06 0.736
2078-09-1.6 2.19e-05 0.880 2.30e-06 0.876
2078-09-7.0 5.66e-07 1.719
2079-08-13.9 1.44e-06 0.888 2.10e-06 0.886
2079-09-9.8 2.36e-07 0.871 1.20e-08 0.719
2079-09-8.5 5.64e-07 0.862 1.30e-06 0.881
2079-09-8.5 4.58e-07 1.716 2.60e-06 1.647
2079-07-30.2 2.95e-07 1.719
2079-09-3.4 6.61e-06 0.888
2080-09-6.2 1.62e-06 0.891 1.30e-06 0.888
2080-09-8.5 4.62e-07 1.712 4.20e-07 1.706
2080-09-7.6 3.28e-07 0.870
2080-09-6.4 2.51e-07 0.862
2080-09-2.8 4.34e-07 0.857
2080-09-6.7 3.32e-08 1.735
2081-09-7.6 2.62e-06 0.895 1.60e-06 0.892
2081-09-10.1 5.07e-07 1.709 4.00e-07 1.703
2081-09-7.0 1.37e-07 0.870
2081-09-3.0 4.14e-07 0.852
2081-08-30.9 2.74e-08 0.720
2081-09-4.1 4.73e-08 1.745
2082-09-9.6 5.60e-07 1.735 1.80e-06 1.321
2082-07-5.0 3.17e-07 0.713
2082-06-29.1 4.71e-08 1.857
2082-09-7.5 2.41e-06 0.898
2082-09-7.3 1.45e-07 0.869
2082-09-3.3 4.58e-07 0.848
2082-09-4.5 1.06e-07 1.745
2083-09-7.2 2.94e-06 0.902 1.90e-06 0.904
2083-09-3.6 4.34e-07 0.847 3.90e-07 1.218
2083-09-9.5 6.60e-07 1.703 5.30e-07 2.433
2083-01-31.8 2.22e-09 0.747
2083-01-30.5 5.14e-09 1.844
2083-06-23.4 5.43e-08 1.857
2083-09-7.2 8.31e-08 0.869
2083-08-31.6 1.96e-07 0.712
2083-09-5.2 9.63e-08 1.756
2083-08-32.3 5.39e-09 1.853
Asteroid 2000 SG344 turns out to be a hard test for our systems. It was discovered on 29 September 2000 by the LINEAR survey at Socorro, New Mexico. The orbital solution used today by NASA JPL is based on observations spanning from 15 May 1999 to 3 October 2000, giving it an observational arc of about 1.4 years, with only a modest number of measurements available for orbit fitting (31 observations). Its orbital period is nearly one year, the inclination relative to the ecliptic is almost zero, and the aphelion lies just beyond Earth’s orbit. These parameters result in very low encounter velocities with Earth (about 1.4 km/s) and a continuum of predicted close approaches after 2070. Because of this unusual dynamical configuration, 2000 SG344 has been the subject of a dedicated impact-hazard study (Chodas & Chesley, 2001).
Sentry-II system indicates 300 possible impacts of this asteroids with the Earth, from which 138 with probabilities higher than $10^{-6}$. NEOForCE system found 1363 possible impacts and 172 with probabilities $10^{-6}$.
5 Conclusions
In this work we present NEOForCE (Near-Earth Object’s Forecast of Collisional Events), a newly developed and fully independent monitoring system for predicting potential asteroid impacts with the Earth. The system is designed to operate autonomously from existing services and will ultimately use orbital solutions provided by the Paris Observatory. A central feature of NEOForCE is its search strategy: rather than testing encounters with the virtual asteroids (VAs) themselves, it identifies epochs when the Earth approaches the curvilinear uncertainty regions of the propagated VAs. This approach, combined with an assessment of the encounter’s maximal impact probability, allows the system to efficiently narrow down possible collisions, and proceed to a targeted search for virtual impactors only when warranted. For the search of virtual impactors we use the semi-linear Partial Banana Mapping method, that can successfully find an impactor, even if the virtual asteroid is at 1 au distance from the Earth.
We validated NEOForCE by comparing its performance with NASA’s Sentry-II system on five asteroids of different orbital and observational characteristics. The results demonstrate that NEOForCE reliably reproduces most of all significant impact solutions reported by Sentry-II above the $10^{-7}$ probability level, confirming the robustness of the method. Moreover, NEOForCE identified additional low-probability impact solutions (in the $10^{-7}$$10^{-6}$ range) that Sentry-II did not report. These tests show that NEOForCE provides a reliable and complementary tool for impact monitoring, strengthening the independence and robustness of global planetary defense capabilities.
Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement #101068341 “NEOForCE”.

### v2

@@
 Robust impact monitoring of near-Earth objects is an essential task of planetary defense.
-Current systems such as NASA’s Sentry-II, University of Pisa’s CLOMON2 and Aegis of ESA have been highly successful, but independent approaches are essential to ensure reliability and to cross-validate predictions of possible impacts, probability and path on Earth.
+Current systems such as NASA’s Sentry-II, NEODyS’s CLOMON2 and Aegis of ESA have been highly successful, but independent approaches are essential to ensure reliability and to cross-validate predictions of possible impacts, probability and path on Earth.
 Aims.
@@
 As other monitoring system NEOForCE samples several thousands of so-called ’virtual asteroids’ from the uncertainty region and integrates their orbits up to 100 years into the future.
-Instead of searching for close approaches of the virtual asteroids themselves with the Earth, our system looks for times when the Earth comes close to the ”realistic” uncertainty regions around them, which are mostly stretched along their osculating orbits.
-For each virtual asteroid we also estimate the maximal impact probability, and only if this value is large enough do we continue to the next step.
+Instead of searching for close approaches of the virtual asteroids themselves with the Earth, our system looks for times when the Earth comes close to the ”realistic” uncertainty regions around them, which are mostly stretched along the osculating ellipses of virtual asteroids.
+For every virtual asteroid and every possible collision time we also estimate the maximal impact probability, and only if this value is large enough ($5\times 10^{-8})$ we continue to the next step.
 In this second step, we compute how the original asteroid orbit should be slightly modified so that the new trajectory leads to an Earth impact, which allows us to confirm the possible collision and estimate the impact probability.
 Results.
-We tested NEOForCE against NASA’s Sentry-II system on five representative asteroids with high impact probability and significant number of possible collisions: 2000 SG344, 2005 QK76, 2008 JL3, 2023 DO, and 2025 JU.
+We tested NEOForCE against NASA’s Sentry-II system on five representative asteroids with high impact probability and significant number of possible collisions: 2000 SG344, 2005 QK76, 2008 JL3, 2023 DO, and 2008 EX5.
 NEOForCE successfully recovered mostly all possible collisions reported by Sentry-II with impact probabilities above $10^{-7}$, demonstrating the robustness of our approach.
@@
 These sets of orbital parameters are in general called virtual asteroids (VAs) (Milani et al., 2000).
-Some of these VAs can collide with the Earth and some pass by safely, which creates an impact probability different from zero or equal to one.
-One of the most natural and general way to estimate the impact probability is the Monte-Carlo approach: sample a large number of VAs from the uncertainty region of orbital parameters according to the probability distribution, propagate their orbit in the future, and last derive impact probability by the fraction of colliding orbits over all all considered in the sample.
-This approach has many advantages: It does not have additional assumptions, and also has a direct formula for estimating the uncertainty of the obtained result: $\sigma_{MC}=\sqrt{\frac{P_{MC}(1-P_{MC})}{n}}$, where $P_{MC}$ is the impact probability and $n$ is the number of considered VAs (Vavilov & Medvedev, 2015).
-However, if the impact probability is small—which is generally the case with PHAs—the number of required VAs to integrate to roughly estimate the impact probability (so that $\sigma_{MC}\lesssim 0.3P_{MC}$) is $10/P_{MC}$.
+Some of these VAs can collide with the Earth and some pass by safely, which allows the impact probability to be in a range between 0 and 1.
+One of the most natural and general way to estimate the impact probability is the Monte-Carlo approach: sample a large number of VAs from the uncertainty region of orbital parameters according to the probability distribution, propagate their orbit in the future, and last derive impact probability by the fraction of colliding orbits over all considered in the sample.
+This approach has many advantages: It does not have additional assumptions, and also has a direct formula for estimating the uncertainty of the obtained result: $\sigma_{MC}=\sqrt{\frac{P_{MC}(1-P_{MC})}{n}}$, where $P_{MC}$ is the impact probability and $n$ is the number of considered VAs (Vavilov and Medvedev, 2015).
+However, if the impact probability is small—which is generally the case with NEAs —the number of required VAs to integrate to roughly estimate the impact probability (so that $\sigma_{MC}\lesssim 0.3P_{MC}$) is $10/P_{MC}$.
 Therefore, this approach is impractical to systematic search for possible impacts with small impact probability values on a large set of of objects.
@@
 The first linear method was introduced by Chodas (1993) who also proposed to use the target plane for the impact analysis.
-The important drawback of this approach is that the asteroid must come close to the Earth (enter its sphere of influence $\approx 929,000$ km), making linearization validity limited.
-Vavilov & Medvedev (2015) proposed the use of a special curvilinear coordinate system related to the asteroid’s osculating orbit to overcome this limitation.
+In addition to the linear assumption (which can be broken by a close encounter) the important drawback of this approach is that the asteroid must come close to the Earth (enter its sphere of influence $\approx 929,000$ km), making linearization validity limited.
+Vavilov and Medvedev (2015) proposed the use of a special curvilinear coordinate system related to the asteroid’s osculating orbit to overcome this limitation.
 This allows us to calculate satisfactorily accurate probabilities of an asteroid collision with the Earth, even if the distance between the Earth and the nominal asteroid at time $t$ exceeds 1 au.
@@
 Because the method constructs the curvilinear uncertainty region of an asteroid then finds the part of this region, which is closest to the Earth and maps it to the target plane.
-However, the linear relation between the errors of orbital parameters is a very important assumption and is limiting the usage of this method to the situation when the two-body formalism is not significantly violated.
+However, the linear relation between the errors of orbital parameters is a very important assumption and is limiting the usage of linear methods to the situation when the two-body formalism is not significantly violated.
 A better compromise between linear methods and a full Monte-Carlo approach was introduced by Milani (1999) as a concept of line of variations (LOV).
-Instead of sampling the full multi-dimensional phase-space, the LOV reduces the problem to a single dimension that follows the direction where uncertainties and nonlinear effects are strongest (Milani et al., 2002, 2005a, 2005b).
+Instead of sampling the full multi-dimensional phase-space, the LOV reduces the problem to a single dimension that follows the direction where uncertainties and nonlinear effects are strongest (Milani et al., 2002, 2005b, 2005a).
 By distributing VAs along this line and interpolating between them, one can efficiently identify possible impacting solutions and trace distinct dynamical pathways.
 This approach makes it possible to detect even very small impact probabilities while keeping the computational effort manageable.
-Because of these advantages, operational monitoring systems at NASA, Sentry [^1] https://cneos.jpl.nasa.gov/sentry/, and CLOMON2222https://newton.spacedys.com/neodys2/index.php?pc=4.1 (operated by the University of Pisa) were both built upon the LOV technique.
+Because of these advantages, operational monitoring systems at NASA, Sentry [^1] https://cneos.jpl.nasa.gov/sentry/, CLOMON2222https://newton.spacedys.com/neodys2/index.php?pc=4.1 (operated by NEODyS) and ESA’s Aegis [^3] https://neo.ssa.esa.int/risk-list were all built upon the LOV technique.
 Recently, NASA introduced the Sentry-II system (Roa et al., 2021), which samples the full confidence region rather than restricting to LOV.
@@
 These methods require several thousands of VAs’ orbits to integrate for the first search of close encounters only, but can detect possible impacts with probabilities of the order of $10^{-7}$ while being much more efficient than the Monte-Carlo one.
-Vavilov & Hestroffer (2025) introduced a semi-linear method for impact probability estimation, as an extension of the linear Partial Banana Mapping method (Vavilov, 2020) built in previous work.
+Vavilov and Hestroffer (2025) introduced a semi-linear method for impact probability estimation, as an extension of the linear Partial Banana Mapping method (Vavilov, 2020) built in previous work.
 This extension (called Impactor PBM or Improved PBM: IPBM) consists of a direct search for a state vector at the epoch of observations, $t_{0}$, that leads to a close encounter at the time of a possible collision, $t$.
@@
 1.1 Concept of the Partial Banana Mapping method
-We briefly outline the concept of the Partial Banana Mapping (PBM) method, which forms the core of our system. A more detailed discussion can be found in Vavilov & Medvedev (2015); Vavilov (2020)
+We briefly outline the concept of the Partial Banana Mapping (PBM) method, which forms the core of our system. A more detailed discussion can be found in Vavilov and Medvedev (2015); Vavilov (2020).
 The key assumption of PBM is that, over a considered time span, the uncertainties of the orbital elements evolve almost linearly from their values at the epoch of observations $t_{0}$.
@@
 To describe this ”banana-shaped” uncertainty region more accurately, we use a special coordinate system based on the osculating elliptical elements.
-In this system, one of the coordinates is the mean anomaly, $M$, and the other two $(\xi,\eta)$ describe small shifts to the side and above/below the orbit (Vavilov & Medvedev, 2015).
+In this system, one of the coordinates is the mean anomaly, $M$, and the other two $(\xi,\eta)$ describe small shifts to the side and above/below the orbit (Vavilov and Medvedev, 2015).
 This makes the uncertainty region easier to describe mathematically.
@@
 2 NEOForCE monitoring system
-Our impact monitoring system Near-Earth Object’s Forecast of Collisional Events uses a state vector (coordinates and velocities) and its 6-dimensional covariance matrix of an asteroid at the average time of observations.
+Our impact monitoring system Near-Earth Object’s Forecast of Collisional Events uses a state vector (coordinates and velocities) and its $6\times 6$ covariance matrix of an asteroid at the average time of observations.
 The covariance matrix $\mathbf{C}_{0}$ gives us an approximation of the uncertainty of the state vector $\mathbf{x}_{0}$ at time $t_{0}$.
-If inside this uncertainty region there is a vector of orbital parameters that leads to a collision with the Earth, then the asteroid could potentially have an impact.
-The goal of a monitoring system is to find all significantly large areas of orbital parameters that lead to collisions.
+If inside this uncertainty region there is a possible initial state vector that leads to a collision with the Earth, then the asteroid could potentially have an impact.
+The goal of a monitoring system is to find all significantly large areas of state vectors that lead to collisions, later referred as Virtual Impactors (VI).
 Our monitoring system consists of several stages.
@@
 The principal vector corresponding to the highest eigenvalue $\lambda_{11}$ determines the LOV.
-We sample $n$ virtual asteroids (VAs) (in general $n=5001$) on this line according to Laplace distribution so that the Laplace integral ($\frac{\sqrt{2}}{2}\exp(-\sqrt{2}|x|)$) between two subsequent points is constant (and equals the integral from the last point to infinity).
+We sample $n$ VAs (in general $n=5001$) on this line according to Laplace distribution so that the Laplace integral ($\frac{\sqrt{2}}{2}\exp(-\sqrt{2}|x|)$) between two subsequent points is constant (and equals the integral from the last point to infinity).
 We use $\sqrt{2}$ to make the dispersion of the distribution equal one.
-Since we consider 5001 VAs the integral between the points equals $b=1/5002$.
+Since we consider $n$ VAs the integral between the points equals $b=1/(n+1)$.
 The VAs are then determined by the following:
-$l_{i}=-\log(1-2ib)/\sqrt{2}*\sqrt{\lambda_{11}};\ i=-2500,2500,$ (2)
+$l_{i}=-\frac{1}{\sqrt{2}}\log(1-2ib)\sqrt{\lambda_{11}};\ \ l_{-i}=-l_{i};\ \ i=0,{n/2}$ (2)
 where $l_{i}$ is the difference along the LOV from the nominal state vector.
@@
 We assign the covariance matrix to each VA by exchanging the largest eigenvalue by $\lambda_{11}^{*}=\max(l_{i}-l_{i-1},l_{i+1}-l_{i})$.
+Thus, a 1-sigma uncertainty ellipsoid around each VA would slightly touch or even include the neighbor VA.
 The covariance matrix determining the uncertainty region for a VA is computed by:
@@
 2.2 Initial check
+Figure 2: Upper chart.
+The histogram of minimal MOID value for asteroids presented on CNEOS web-site as ones that have non-zero impact probabilities with the Earth on September 9, 2025.
+Bottom chart.
+The cumulative histogram of minimal MOID for all NEOs.
+The vertical red line is our empirical 0.02 au threshold.
 Instead of running NEOForCE system for all near-Earth asteroids we first fulfill the initial check whether this asteroid possibly can come close to the Earth.
-We select 17 VAs along LOV (on 5-sigmasigma interval equidistantly), integrate their orbits into the future and compute minimal orbit intersection distance (MOID) between their orbits and the Earth’s orbit every 15 days.
-If the minimal MOID among all of them is greater than a certain threshold (0.002 au), we don’t run the computation of the impact probability at all.
+We select 100 VAs along LOV (on 5-$\sigma$ interval equidistantly), integrate their orbits into the future and compute minimal orbit intersection distance (MOID) between their orbits and the Earth’s orbit every 100 days.
+If the minimal MOID among all of them is greater than a certain threshold (0.02 au), we don’t run the computation of the impact probability at all.
+Fig.
+2 presents this minimal MOID value for all the asteroids, which are in the impact risk table on CNEOS web-site (on September 9, 2025), as well as the minimal MOID for all NEOs.
+As one can see our threshold 0.02 au includes all currently known NEAs with non-zero impact probability.
+At the same time it allows us to limit the number of launches of the monitoring system to approximately 15,000 instead of 40,000.
+This initial check requires less than one minute of computations for one asteroid on one thread of a processor like Intel(R) Xeon(R) Gold 6230 2.10GHz.
 2.3 Close encounter precomputation
@@
 We integrate the orbit of a VA with variational equations towards the future until the final time (in general, 100 years).
-On each integration step we compute the osculating orbit of a VA and then obtain the minimal distance between the position of the Earth and this osculating ellipse (or parabola/hyperbola) numerically.
+On each integration step we compute the osculating orbit of a VA and then obtain the minimal distance between the position of the Earth at this time and the points of this osculating ellipse (or parabola/hyperbola) numerically.
 We then find the local minima in these distances indicating the time of the local minima $t$, the state vector of the VA, $\mathbf{x}^{*}_{t}$, and the Earth and the partial derivative matrix of VA’s state vector at $t$ over original time $t_{0}$, $\mathbf{\Phi}(t_{0},t)$.
@@
 1) and compute the probability by the target plane method (Chodas, 1993).
-To account for the distance from the VA to the nominal orbit we then multiply this value by $\exp(-\frac{1}{2}\frac{\delta^{2}}{\lambda_{11}})$, where $\delta$ is the deviation in terms of LOV of this VA from zero (i.e.
-the distance on LOV of the considered VA from the nominal asteroid).
+To account for the distance from the VA to the nominal orbit we then multiply this value by $\exp(-\frac{1}{2}\frac{l_{i}^{2}}{\lambda_{11}})$, where $l_{i}$ is from Eq.
+2.
 This probability tells us what is the maximal possible value of the impact probability with the Earth selected asteroid can have at approximately time $t$ around the vicinity of this VA.
@@
 Again, we follow the idea of Milani et al.
-(2005a) in CLOMON2 system, that in order to confirm a possible collision, one must explicitly find a virtual impactor (a state vector at initial epoch, $t_{0}$, which actually leads to a collision).
+(2005b) in CLOMON2 system, that in order to confirm a possible collision, one must explicitly find a virtual impactor (a state vector at initial epoch, $t_{0}$, which actually leads to a collision).
 Therefore, in this stage we try to find the virtual impactor (VI) around each representative found in the previous step.
-For this we use a semi-linear approach extension of the PBM, briefly described here, for details see Vavilov & Hestroffer (2025).
+For this we use a semi-linear approach extension of the PBM, briefly described here, for details see Vavilov and Hestroffer (2025).
 The objective is to find the initial condition at epoch $t_{0}$ of the asteroid, bringing it at time $t$ to the closest point B in Fig. 1. The linear approximation reads as follows:
-$\bm{\xi}_{*}-\bm{\xi}=\left[\frac{\partial\xi}{\partial x_{0}}\right](\bm{x}_{0*}-\bm{x}_{0}),$ (6)
+$\bm{\xi}_{*}-\bm{\xi}=\left[\frac{\partial\bm{\xi}}{\partial\bm{x_{0}}}\right](\bm{x}_{0*}-\bm{x}_{0}),$ (6)
 where $\bm{\xi}$ and $\bm{\xi}_{*}$ are vectors in $(\xi,\eta,M)$ coordinate system at time $t$ of the asteroid (point A) and VA closest to the Earth on the main axis of curvilinear uncertainty ellipsoid (point B), correspondingly.
@@
 $(\bm{x}_{0*}-\bm{x}_{0})$ is the vector we are looking for: how to change the state vector of the selected VA on the previous stage so that after orbit integration it goes to point B at time $t$.
-In practice, for NEOForCE system we also consider: $\frac{3}{4}(\bm{x}_{0*}-\bm{x}_{0})$, $\frac{1}{2}(\bm{x}_{0*}-\bm{x}_{0})$ and state vectors.
+In practice, for NEOForCE system we also use these two additional state vectors: $\frac{3}{4}(\bm{x}_{0*}-\bm{x}_{0})+\bm{x}_{0}$ and $\frac{1}{2}(\bm{x}_{0*}-\bm{x}_{0})+\bm{x}_{0}$.
+These state vectors are closer to $\bm{x}_{0}$ than $\bm{x}_{0*}$.
 We integrate these orbits until time $t$ plus 0.5 days.
@@
 This newly selected VA is closer to the Earth at approximately time $t$ than the original one, so we repeat the procedure using this VA as the new reference (starting from computing $\bm{\xi}$, $\bm{\xi}_{*}$ vectors and matrix $\left[\frac{\partial\xi}{\partial x_{0}}\right]$).
-In this way, the method works as an iterative refinement process: at each step the candidate VA is moved closer to the Earth at $\approx t$ until one of three conditions is met: a collision is found, the solution drifts outside the 9-$\sigma$ uncertainty ellipsoid, or the maximum number of 10 iterations is reached.
-In an extreme case where the stretching of the uncertainty region is very high and mismatches the linear estimation, the closest VA between these 3 cases can be further away than the original one.
-In this case we select $\frac{1}{6}(\bm{x}_{0*}-\bm{x}_{0})$ virtual asteroid and proceed again.
-Once the VA entered the sphere of influence (approximately 929,000 km) we launch the target plane analysis (see Section 3).
+In this way, the method works as an iterative refinement process: at each step the candidate VA is moved closer to the Earth at $\approx t$ until one of four conditions is met:
+  1. (i)
+    a collision is found: we compute the probability by the target plane method (see Section 3) ;
+  2. (ii)
+    the solution drifts outside the 9-$\sigma$ uncertainty ellipsoid: we stop the procedure
+  3. (iii)
+    the maximum number of 10 iterations is reached: we stop the procedure;
+  4. (iv)
+    the asteroid enters the sphere of influence with the Earth — we run the target plane analysis (see Section 3) to search for a VI.
+By taking two additional state vectors, which are closer to original one, instead of simply using $\bm{x}_{0*}$ we slightly mitigate some convergence problems that can happen after multiple close encounters.
+In an extreme case where the stretching of the uncertainty region is very high and mismatches the linear estimation, the closest VA between these 3 cases can be further away from the Earth than the original one.
+In this case we select $\frac{1}{6}(\bm{x}_{0*}-\bm{x}_{0})+\bm{x}_{0}$ VA and proceed again.
 If virtual impactor is found, $\bm{x}_{0*}$, we estimate its impact probability by the target plane method and multiply this value by
@@
 The red denotes the VI.
-On the bottom chart one can see the nominal position of the asteroid at time of a possible collision, $t$ (point A), the osculating orbit of the nominal asteroid (black part of an ellipse), the Earth, the position of the VA at time $t$ (green dot), the uncertainty region assigned for this VA at time $t$ (green curved ellipse).
+On the bottom chart one can see the nominal position of the asteroid at time slightly before a possible collision, $t$ (point A), the osculating orbit of the nominal asteroid (black part of an ellipse), the Earth, the position of the VA (green dot), the uncertainty region assigned for this VA at time slightly before $t$ (green curved ellipse).
 The red dot (point B) represents the position of the VI.
@@
 The Earth collides with the red dot.
-A schematic illustration of the work of our system is presented in Fig. 3.
 2.6 Combination of similar impactors
@@
 We publish a possible collision if the impact probability is higher than $10^{-10}$.
-The reason why this threshold is significantly higher than $5\cdot 10^{-8}$ threshold for the maximal impact probability is that if the maximal impact probability is too small it means the uncertainty of the asteroid around this area is too large.
+The reason why this threshold is significantly higher than $5\times 10^{-8}$ threshold for the maximal impact probability is that if the maximal impact probability is too small it means the uncertainty of the asteroid around this area is too large.
 On the other hand, if the maximal impact probability is not too small but the actual impact probability of a VI is very low, this means that the uncertainty region only grazes the Earth rather than passing through its center.
 In such cases, even small changes in the nominal orbital solution can lead to large fluctuations in the estimated impact probability, which is why these cases are interesting.
+2.7 Short summary of the method
+Below we provide a concise overview of the NEOForCE impact-monitoring algorithm. A schematic illustration of the work is shown in Fig. 4.
+  1. 1.
+    Sampling. We create many virtual asteroids (VAs) along the Line of Variations (LOV) at the mean observation epoch (Section 2.1).
+  2. 2.
+    Initial screening. We make a first quick test to see whether the asteroid can ever come close to the Earth. For this, we compute the minimum MOID for a small group of VAs (Section 2.2).
+  3. 3.
+    Propagation and close-approach detection.
+We integrate the orbit of each VA into the future.
+We record every moment when the Earth comes close to the osculating orbit of a VA (osculating ellipse), together with the time and distance (Section 2.3).
+  4. 4.
+    Curvilinear-region analysis.
+For every case where the distance is smaller than 0.05 au, we build the curvilinear uncertainty region of the VA and compute its distance to the Earth.
+We also estimate the maximum impact probability that this VA could have (Section 2.3).
+  5. 5.
+    Selection of representative scenarios.
+From the previous step, we select only the most important situations: VAs that are neighbors along LOV and bracket the Earth at approximately the same time, have small distances between their uncertainty regions and the Earth, and show a noticeable impact probability (Section 2.4).
+  6. 6.
+    Search for virtual impactors.
+For each selected case, we try to improve the VA’s state vector so that the VA hits the Earth.
+We do this step by step, coming closer to an impact solution.
+When the VA enters the Earth’s sphere of influence at the probable collision time, we use the target-plane analysis (Sections 2.5 and Section 3).
+  7. 7.
+    Combine similar impactors.
+If two virtual impactors stay very close to each other on their way toward the Earth, we treat them as the same solution and keep the one with the higher impact probability (Section 2.6).
 3 Target plane analysis
@@
 An asteroid will collide with the Earth if the distance from the Earth’s center is smaller than its radius $R_{\oplus}$ (see Fig.
-4).
+5).
 Because of gravitational focusing, however, the asteroid’s trajectory inside the Earth’s sphere of influence is hyperbolic, so the effective cross-section is larger.
 The minimum distance $q$ is given by
-$q=\frac{b}{\sqrt{1+\frac{v_{s}^{2}}{v_{rel}^{2}}}},$ (11)
-where $v_{s}\approx 11.186\ \mathrm{km\ s^{-1}}$ is the Earth’s surface escape velocity and $b$ is the impact parameter (the distance on the target plane between the projections of the Earth’s and asteroid’s centers).
+$q=\frac{b}{\sqrt{1+\frac{v_{s}^{2}}{v_{\infty}^{2}}}},$ (11)
+where $v_{s}\approx 11.186\ \mathrm{km\ s^{-1}}$ is the Earth’s surface escape velocity, $v_{\inf}$ is the unperturbed geocentric velocity and $b$ is the impact parameter (the distance on the target plane between the projections of the Earth’s and asteroid’s centers).
 Figure 5: Schematic illustration of the target plane method.
@@
 To address this, we adopted a modified approach similar to Milani et al.
-(2005a).
+(2005b).
 We extend the integration until the epoch of closest approach and construct a ”modified target plane (MTP)” perpendicular to the relative velocity at that instant.
@@
 We therefore map $\mathbf{C}_{xyz\dot{x}\dot{y}\dot{z}}$ onto TP1, rotate by $\alpha$, and scale by $sc$ to obtain the $2\times 2$ matrix $\mathbf{C}_{TP^{*}}$.
-3.2.1 Search for virtual impactor
+3.3 Search for virtual impactor
 To confirm whether a potential encounter corresponds to a genuine collision, we follow the method of Milani et al.
-(2005a).
+(2005b).
 The 6D state vector at $t$ relative to the nominal state vector of the asteroid can be partitioned into $\bm{r},\bm{s}]$, where $\bm{r}=(\xi_{TP},\zeta_{TP})$ are target-plane coordinates and $\bm{s}$ is the complementary 4D vector.
@@
 $\mathbf{C}=\left(\begin{array}[]{cc}\mathbf{C_{r}}&\mathbf{C_{rs}}\\ \mathbf{C_{sr}}&\mathbf{C_{s}}\end{array}\right).$ (15)
-According to (Milani et al., 2005a) , vector $\bm{s_{*}}$ can be found from
+According to (Milani et al., 2005b) , vector $\bm{s_{*}}$ can be found from
 $\bm{s_{*}}=\bm{s}+\mathbf{C_{r}}^{-1}\mathbf{C_{sr}}(\bm{r}-\bm{r_{*}}).$ (16)
@@
 As in Section 2.5 we also consider $\frac{3}{4}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$ and $\frac{1}{2}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$ and select the closest one to the Earth at time $t$.
-If all these three are more distant than $\bm{x}_{0*}$ at time $t$ we take $\frac{1}{6}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$.
+If all these three are more distant than $\bm{x}_{0*}$ at time $t$ we take $\frac{1}{10}(\bm{\tilde{x}}_{0*}-\bm{x}_{0*})$.
 If the state vector $\bm{\tilde{x}}_{0*}$ does not lead to a collision approximately at time $t$, this process needs to be iterated taking $\bm{\tilde{x}}_{0*}$ as $\bm{{x}}_{0*}$ and using Eqs.
@@
 4 Practical realization and discussion
-In this version of the system we used Everhart (1974) integrator of 15-th order.
-In the model we included gravitational influence of 8 major planets (and Pluto for historical reasons).
+In this version of the system we used collocational integrator Lobbie of 14-th order (Avdyushev, 2022) for integrating the orbits.
+In the model we included gravitational influence of 8 major planets (and dwarf planet Pluto, for historical reasons), as well as post-Newtonian formalism for the Sun only.
 We also considered the Earth and the Moon separately.
 The orbits of asteroids and their covariance matrices are taken from DynAstVO database [^4] https://epn.imcce.fr/dynastvo.html Desmars et al.
-(2017).
+(2017) for the automated impact monitoring.
 4.1 Completeness problem
 The completeness of our system is determined by the threshold of impact probability that can be detected certainly.
-Following (Milani et al., 2005a) one can estimate the generic completion level: the highest impact probability that can be missed by our monitoring system if the uncertainty trail on the target plane is fully linear.
+Following (Milani et al., 2005b) one can estimate the generic completion level: the highest impact probability that can be missed by our monitoring system if the uncertainty trail on the target plane is fully linear.
 Under these assumptions, a possible collision can be missed if both representatives found in Section 2.4 are further than 1 au from the Earth.
@@
 This is slightly higher than if we sampled VAs along the LOV according to Gaussian distribution.
-In this case, the probability would be $2\cdot 10^{-4}$.
+In this case, the probability would be $2\times 10^{-4}$.
 Our objective is hence the following: If we took a sample of 5001 VAs according to the Gaussian distribution, the latest VAs would be only 3.54 sigmas away from the nominal state vector ($l_{2500}=3.54$ from equation 2).
@@
 4.2 Comparison to Sentry
-To evaluate the robustness of our system, we performed detailed comparisons using five representative asteroids: 2005 QK76, 2008 JL3, 2023 DO, 2025 JU, and 2000 SG344.
+To evaluate the robustness of our system, we performed detailed comparisons using five representative asteroids: 2005 QK76, 2008 JL3, 2023 DO, 2008 EX5, and 2000 SG344.
 We selected these asteroids from the web-page Sentry of Central for Near-Earth Objects Study (CNEOS) [^5] https://cneos.jpl.nasa.gov/sentry.
@@
 The asteroids have different observational arcs (from 1.7 days to 1.4 years), and different number of observations (from 14 to 52).
-For each case, we adopted the same orbital solutions as those used by NASA’s Sentry-II system, retrieved via the NASA SSD API [^5] https://ssd-api.jpl.nasa.gov on September 9, 2025, so that the results would be directly comparable.
+For each case, we adopted the same orbital solutions and the same covariance matrices as those used by NASA’s Sentry-II system, retrieved via the NASA SSD API [^6] https://ssd-api.jpl.nasa.gov on September 9, 2025, so that the results would be directly comparable.
 The outcomes of our system and Sentry-II are presented in the tables below.
@@
 We attempted to match virtual impactors (VIs) such that, for collisions occurring at similar epochs, both the probability and the distance metric $\sigma_{VI}$ are comparable.
-In our approach, $\sigma_{VI}$ quantifies how far a VI lies from the nominal orbit in terms of standard deviations and is defined as
+$\sigma_{VI}$ quantifies how far a VI lies from the nominal orbit in terms of standard deviations and is defined as
 $\sigma_{VI}^{2}=[\bm{x}_{0*}-\bm{x_{0}}]^{\mathrm{T}}\mathbf{C}_{0}^{-1}[\bm{x}_{0*}-\bm{x_{0}}],$ (19)
 where $\bm{x}_{0*}$ is the state vector of the VI, $\bm{x}_{0}$ is the nominal state vector, and $\mathbf{C}_{0}$ is the covariance matrix.
-Sentry-II estimates $\sigma{VI}$ using the root-mean-square residuals of the observational fit for the impacting solution.
 It is important to note that similar $\sigma_{VI}$ values do not necessarily imply that the same VI was found: many virtual asteroids can lie on the same 5D ellipsoidal surface defined by a given $\sigma_{VI}$.
 Conversely, two very similar VIs may have different $\sigma_{VI}$ values, for example, a small shift along the direction of the smallest eigenvector of $\mathbf{C}_{0}$ can leave the overall trajectory nearly unchanged while significantly altering $\sigma_{VI}$.
-Table 1: Comparison table for asteroid 2025 JU.
+Table 1: Comparison table for asteroid 2008 EX5.
 Date $P_{NEOForCE}$ $\sigma_{VI}$ $P_{NASA}$ $\sigma_{VI_{NASA}}$
-2074-10-2.3 4.10e-05 1.091 4.30e-05 1.539
-2074-10-2.3 1.49e-05 0.998 1.50e-05 1.500
-2076-10-1.9 1.71e-07 3.268 1.70e-07 3.855
-2079-10-2.5 1.91e-08 0.949 2.00e-08 1.362
-2079-10-2.5 8.04e-08 0.947 8.00e-08 1.193
-2079-10-2.5 6.25e-07 0.742 6.80e-07 0.956
-2079-10-2.5 5.68e-07 0.741 5.50e-07 0.937
-2091-10-2.4 5.50e-10 1.182
-2091-10-2.5 3.56e-09 1.830
-2103-10-3.3 5.25e-09 3.011
-2108-10-2.6 2.21e-08 1.164 2.90e-08 1.929
-2108-10-2.5 7.23e-08 1.170 6.70e-08 1.471
-2108-10-2.6 3.13e-07 1.175 3.10e-07 1.675
-4.2.1 Asteroid 2025 JU
-Asteroid 2025 JU is newly discovered, with 38 observations spanned over 35 days.
-Its cumulative impact probability is about $6\times 10^{-5}$, not very high and it also has only a few significant possible impacts.
-Table 1 shows an iconic agreement between two systems with Sentry loosing three possible collisions with a probability less than $5.3\times 10^{-9}$ (way below the mentioned completeness level).
+2056-10-8.9 7.04e-08 0.504 7.00e-08 1.028
+2059-10-9.5 5.38e-08 0.505
+2061-10-9.0 3.00e-08 0.501 3.00e-08 1.400
+2062-10-9.5 6.64e-07 0.562 7.80e-07 0.604
+2063-10-9.5 5.09e-08 0.498 4.80e-08 0.618
+2065-10-9.0 4.41e-08 0.506 4.70e-09 0.564
+2065-10-09.09 4.50e-08 1.085
+2067-10-9.6 4.11e-08 0.503
+2067-10-9.8 9.43e-08 0.710
+2068-10-9.0 5.09e-07 0.551 6.90e-07 0.853
+2068-10-08.85 3.50e-10 2.175
+2069-10-9.1 1.02e-07 0.255 1.00e-07 1.324
+2070-10-09.39 6.20e-08 0.649
+2071-10-9.7 2.10e-06 0.371 2.00e-06 0.644
+2071-10-09.63 3.40e-08 1.238
+2072-10-9.0 1.94e-05 0.919 2.70e-05 0.922
+2072-10-8.9 1.37e-06 1.533 1.40e-06 1.685
+2074-10-9.5 4.62e-07 0.481 4.80e-07 0.592
+2074-10-9.4 1.26e-07 1.615 1.30e-07 1.707
+2075-10-9.8 2.19e-08 0.917 6.70e-10 1.449
+2076-10-8.9 6.60e-08 0.394 6.30e-08 0.447
+2076-10-8.9 2.52e-08 0.420
+2077-10-9.1 3.11e-07 0.542 3.50e-07 0.699
+2079-10-9.8 3.74e-08 0.266 3.80e-07 0.286
+2079-10-9.6 3.45e-08 0.578 3.40e-08 1.161
+2080-10-9.0 5.28e-08 0.715
+2081-10-9.1 8.56e-08 1.579
+2082-10-9.3 9.08e-08 1.620 9.30e-08 2.029
+2082-10-9.3 1.80e-10 0.576
+2083-10-9.7 1.89e-05 0.844 2.00e-05 1.002
+2083-10-09.66 1.30e-08 0.617
+2084-10-8.9 5.56e-08 1.582 5.40e-08 1.619
+2086-10-9.5 4.42e-10 0.257
+2088-10-09.02 5.40e-09 1.445
+2090-10-9.5 5.44e-08 0.030 5.80e-08 1.282
+2093-10-09.25 2.30e-09 0.045
+4.2.1 Asteroid 2008 EX5
+Asteroid 2008 EX5 is newly discovered, with 61 observations spanned over 32 days.
+Its cumulative impact probability is about $4.5\times 10^{-5}$, not very high and it also has only a few significant possible impacts.
+Table 1 shows an agreement between two systems with Sentry loosing 8 possible collisions with a probability less than $1\times 10^{-7}$ (below the mentioned completeness level) and NEOForCE looses 7 possible collisions with probabilities less than $5\times 10^{-8}$, below the threshold for maximal impact probability.
 Table 2: Comparison table for asteroid 2005 QK76.
@@
 Such a short observation arc makes its orbit highly uncertain and a good test for comparison.
-Sentry-II does not use the covariance matrix only the observations itself, while NEOForCE uses the covariance matrix in coordinates and velocities at the mean epoch of observations.
 As one can see from Table 2, there is no significant difference in the impact probabilities of the possible collisions found.
@@
 Discovered on May 2, 2008, it has 44 possible collision solutions listed in Sentry-II.
-A comparison of our system with Sentry-II (Table NEOForCE: Near-Earth Objects’ Forecast of Collisional Events) shows that NEOForCE identified 87 possible collisions but did not recover 3 mentioned by Sentry-II, all with impact probabilities below $4.6\times 10^{-9}$ — way below our completeness threshold.
+A comparison of our system with Sentry-II (Table 4.2.5) shows that NEOForCE identified 87 possible collisions but did not recover 3 mentioned by Sentry-II, all with impact probabilities below $4.6\times 10^{-9}$ — way below our completeness threshold.
 In contrast, Sentry-II missed 46 possible collisions, including one with the highest reported impact probability of $1.26\times 10^{-6}$ on April 28, 2109, as well as three others with probabilities exceeding $10^{-7}$.
@@
 Its orbital period is nearly one year, the inclination relative to the ecliptic is almost zero, and the aphelion lies just beyond Earth’s orbit.
-These parameters result in very low encounter velocities with Earth (about 1.4 km/s) and a continuum of predicted close approaches after 2070.
-Because of this unusual dynamical configuration, 2000 SG344 has been the subject of a dedicated impact-hazard study (Chodas & Chesley, 2001).
+These parameters result in very low encounter velocities with Earth (about 1.4 km/s) and a continuum of predicted close approaches after several resonance returns after 2070.
+Because of this unusual dynamical configuration, 2000 SG344 has been the subject of a dedicated impact-hazard study (Chodas and Chesley, 2001).
 Sentry-II system indicates 300 possible impacts of this asteroids with the Earth, from which 138 with probabilities higher than $10^{-6}$.
 NEOForCE system found 1363 possible impacts and 172 with probabilities $10^{-6}$.
+For the reasons discussed above, the computation time for asteroid 2000 SG344 was significantly higher (more then 6x longer than for 2008 JL3).
+Table 6 reports the total computation time, obtained in parallel mode using two threads on an Intel(R) Xeon(R) Gold 6230 CPU at 2.10 GHz, together with the number of situations in which the virtual impactor (VI) search procedure was triggered.
+Table 5: Comparison table for asteroid 2008 JL3
+Date $P_{NEOForCE}$ $\sigma_{VI}$ $P_{NASA}$ $\sigma_{VI_{NASA}}$
+2027-04-31.3 1.39e-04 0.111 1.50e-04 0.120
+2030-04-30.8 3.00e-08 4.242 3.00e-08 4.286
+2034-04-30.5 1.67e-07 0.105 1.60e-07 0.110
+2043-05-01.05 4.60e-09 0.117
+2047-05-01.52 3.20e-09 0.116
+2048-04-30.5 1.32e-08 0.115
+2051-04-31.4 7.34e-08 0.104
+2061-04-31.1 5.74e-09 0.115
+2063-05-2.0 2.28e-07 0.092 2.10e-07 0.096
+2065-04-27.7 8.02e-08 0.911 7.90e-08 0.920
+2071-04-30.3 2.63e-06 1.022 2.60e-06 1.032
+2072-04-30.5 2.03e-08 0.117
+2075-05-2.1 8.78e-08 1.598 1.00e-07 1.614
+2075-05-2.0 9.28e-08 0.081
+2077-04-28.5 9.27e-09 2.222
+2080-04-26.93 2.40e-09 0.193
+2081-04-31.5 1.39e-08 0.013
+2082-04-27.4 4.08e-09 0.084 4.60e-09 0.153
+2083-05-1.8 1.60e-08 0.095
+2087-04-31.2 2.03e-08 2.145
+2090-05-2.2 7.96e-08 0.099 7.70e-08 0.098
+2090-05-1.6 1.25e-09 0.158 5.80e-09 0.167
+2090-05-1.6 3.11e-07 0.193 2.80e-07 0.195
+2090-04-31.5 5.40e-07 0.195 5.50e-07 0.197
+2090-04-31.2 6.47e-07 0.229 5.80e-07 0.230
+2090-04-31.5 5.52e-07 1.958 5.30e-07 1.978
+2090-04-31.1 2.27e-08 0.188
+2090-05-1.5 1.14e-08 2.291
+2090-05-1.8 7.40e-08 2.792
+2090-05-1.9 4.76e-08 3.011
+2091-04-28.8 9.86e-09 0.919 9.10e-09 0.923
+2092-04-28.9 2.72e-08 1.041 3.10e-08 1.052
+2092-04-28.5 2.75e-08 0.189
+2093-05-1.6 1.32e-08 0.166
+2095-04-30.8 1.71e-08 0.085 1.60e-08 0.092
+2095-04-31.4 3.21e-08 0.139 2.70e-08 0.139
+2098-04-31.0 1.45e-08 0.190
+2099-04-31.0 6.33e-09 2.141
+2099-04-30.8 1.53e-08 0.084
+2099-04-30.9 1.53e-08 1.023
+2100-04-27.8 1.87e-07 1.267 1.80e-07 1.283
+2100-04-27.9 2.63e-06 0.264 2.50e-06 0.275
+2100-04-27.9 8.81e-07 0.174 8.60e-07 0.176
+2100-04-27.8 1.25e-07 0.131 1.10e-07 0.134
+2100-04-27.4 9.47e-10 2.651
+2100-04-27.8 1.08e-08 2.297
+2102-05-1.8 1.51e-08 0.231 3.40e-09 2.267
+2103-04-28.3 4.90e-08 2.214
+2103-04-28.5 1.86e-09 0.096
+2109-04-27.9 2.55e-07 0.594 2.30e-07 0.601
+2109-04-28.1 6.88e-08 1.089 6.50e-08 1.103
+2109-04-28.1 4.91e-09 0.660
+2109-04-28.0 1.26e-06 0.831
+2109-04-28.2 8.90e-08 1.240
+2110-05-2.9 4.12e-08 2.835 3.90e-08 2.864
+2110-05-2.5 2.61e-08 2.055
+2110-05-2.4 3.34e-08 1.925
+2110-05-2.4 2.89e-08 1.463
+2110-05-2.4 2.01e-09 0.190
+2112-04-28.0 1.93e-07 0.210 1.80e-07 0.214
+2112-04-27.9 3.18e-08 0.221 3.90e-10 0.149
+2112-04-28.4 5.23e-07 0.906 4.90e-07 0.906
+2112-04-28.3 2.49e-07 2.547 2.50e-07 2.573
+2112-04-28.4 1.55e-07 0.903
+2113-04-31.4 4.56e-07 0.115 4.50e-07 0.117
+2113-05-1.8 4.37e-08 0.027 9.30e-09 0.054
+2113-05-2.2 8.91e-10 3.718
+2113-05-1.8 1.68e-07 2.495
+2113-05-2.2 2.41e-08 0.005
+2113-05-1.9 7.93e-09 0.147
+2114-04-31.5 9.87e-09 0.084 9.50e-09 0.116
+2114-05-1.6 1.54e-09 2.235
+2115-04-28.5 3.32e-09 0.181
+2115-04-28.4 1.60e-08 0.980
+2116-04-30.1 1.39e-06 1.324 1.30e-06 1.338
+2116-04-30.4 2.92e-07 0.025 2.60e-07 0.035
+2117-05-2.2 1.85e-07 0.060 1.70e-07 0.061
+2117-04-31.3 6.86e-10 0.188
+2119-05-3.5 3.68e-08 0.234 3.50e-08 0.236
+2119-05-3.5 9.36e-07 0.385 8.20e-07 0.389
+2119-05-3.8 1.65e-10 0.106
+2119-05-3.6 1.05e-07 2.012
+2120-04-29.0 7.36e-10 2.812
+2121-04-28.2 3.90e-08 0.178
+2121-04-28.0 4.69e-09 0.192
+2122-05-3.6 2.11e-08 0.091 2.90e-07 0.238
+2122-05-3.1 2.18e-06 0.362 2.00e-06 0.366
+2122-05-3.5 5.13e-07 2.030 5.10e-07 2.051
+2123-04-30.7 9.69e-09 0.084
+2124-04-27.9 1.47e-09 0.163
+Table 6: Computation times and the number of VI searches for 5 considered asteroids.
+Asteroid $N_{VI}$ $t_{comp}$, s
+2008 EX5 965 2738.8
+2005 QK76 555 2627.7
+2023 DO 2118 6216.5
+2008 JL3 984 2361.9
+2000 SG344 8530 15524.7
+’Asteroid’ is the asteroid designation, $N_{VI}$ is the number of situations in which the virtual impactor (VI) search procedure was triggered, $t_{comp}$ is the total computation time, obtained in parallel mode using two threads on an Intel(R) Xeon(R) Gold 6230 CPU at 2.10 GHz.
+A summary of the comparison for all five considered asteroids with NASA is shown in Fig.
+6.
+For asteroids 2008 EX5, 2005 QK76, 2023 DO, and 2008 JL3, we find good agreement between the two systems, with impact probability values consistent down to the $10^{-7}$ level, except for a few missing collision cases.
+For asteroid 2000 SG344, the agreement is noticeably worse and holds only down to about the $10^{-6}$ level.
+In this case, several potential collisions at the $10^{-5}$ probability level are missed by the NASA system.
+This behavior is likely caused by the very small relative velocity during close encounters, which reduces the completeness of the monitoring system (see Sect.
+4.1).
+Figure 6: Comparison between the NEOForCE impact monitoring system and Sentry-II.This figure summarizes the results from Tables 1–5.
+We show only impact probabilities larger than $10^{-9}$.
+A point located on the horizontal line at $P_{NASA}=10^{-9}$ indicates a potential impact identified by NEOForCE but not detected by Sentry-II.
+Conversely, a point on the vertical line indicates an impact detected by Sentry-II but not found by NEOForCE.
 5 Conclusions